b) 3 g of dihydrogen react with 29.0 g of dioxygen to yield water as follows.

O2 (g) + 2H2 (g) -----------------? 2H2O (g)

1. Which is limiting reactant?
2. Calculate the maximum amount of H2O that can be formed.
3. Calculate the amount of one of the reactants which remains unreacted.

Grade:11

1 Answers

Umakant biswal
7 years ago
@ shaik 
here the equation is balanced , so, we need to calculate no of moles frst , 
h2 = 3/2 = 1.5 moles 
o2 = 29/32 = 0.90 moles 
then just divide the no of moles with the coefficient of balanced equation 
h2 = 1.5 / 2 = 0.75 
o2 = 0.90/ 1= 0.90 
so, here H2 will be the limiting reagent , 
HOPE IT CLEARS